OpenStudy (anonymous):

Reggie is 72 inches tall and weighs 130 pounds. Use the formula below to answer the question about his weight. BMI = Weight (in pounds) ÷ [Height (in inches)]² x 703 Which BMI category does Reggie fall into?

OpenStudy (anonymous):

BMI = 130 / (72^2) x 703 BMI = 130/5184 x 703 BMI = . 025 x 703 BMI = 17.58 categories.... underweight < 18.5 normal weight 18.5 - 24.9 overweight 25 - 29.9 obesity 30 or greater Reggie falls in the underweight category.